OpenStudy (anonymous):

A = 1/2 h(B + b). Solved for b, this formula becomes b =

OpenStudy (anonymous):

it is \[A=\frac{1}{2}h(B+b)\]?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

1) multiply both sides by 2 to get rid of the fraction what do youg get?

OpenStudy (anonymous):

a*2=h(B+b)?

OpenStudy (anonymous):

yeah or \[2A=h(B+b)\]

OpenStudy (anonymous):

now divide both sides by \(h\)

OpenStudy (anonymous):

by which i mean just write it you don't actually perform a division just write \[\frac{2A}{h}=B+b\]

OpenStudy (anonymous):

then to get \(b\) by itself, subtract \(B\) from both sides